Please help me solve my assignment. Thank you.
Points N, R and T lie in a straight line. R lies between N and T. If RT = 4/3 NT
and RT = 100, find NT.

Answer by checkley75(3666)

RT CANNOT BE GREATER [4/3=1.333] THAN NT IF R IS BETWEEN N&T
PROOF:
RT=100 & RT=(4/3)NT THEN
100=(4/3)NT
NT=100/(4/3)
NT=100(3/4)
NT=300/4
NT=75
THE ENTIRE LENGTH OF THE LINE (NT=75) CANNOT BE LONGER THAN THE PART (RT=100)